OpenStudy (anonymous):

Square root Lnx/Ln square root of x= 2 How do I solve this?

OpenStudy (anonymous):

Square and exponentiate.

OpenStudy (anonymous):

Square what?

OpenStudy (anonymous):

Everything

OpenStudy (anonymous):

No idea?

OpenStudy (anonymous):

I still have no idea how to figure it out.. sorry logs and natural logs confuse me.

OpenStudy (nikvist):

do you think \[\frac{\sqrt{\ln{x}}}{\ln{\sqrt{x}}}=2\]

OpenStudy (nikvist):

\[\frac{\sqrt{\ln{x}}}{\ln{\sqrt{x}}}=2\quad,\quad x>1\] \[\frac{\sqrt{\ln{x}}}{0.5\cdot\ln{x}}=2\] \[\frac{\sqrt{\ln{x}}}{\ln{x}}=1\quad,\quad y=\sqrt{\ln{x}}\] \[\frac{y}{y^2}=1\quad\Rightarrow\quad y=1\quad\Rightarrow\quad x=e\]

OpenStudy (anonymous):

Squaring gives ln x = (ln x)^2 true for x = 1 or e We have to discard 1 because division by 0.